# Rad4XCNN: a new agnostic method for post-hoc global explanation of CNN-derived features by means of radiomics

### v1

@@ 6 Conclusion
Our findings highlight significant differences between radiomic, CNN-derived, and ViT-derived features in terms of accuracy and explainability. While deep architectures, particularly CNNs, demonstrate superior performance in terms of accuracy, radiomic features offer clearer and more comparable explanations due to their well-defined significance. Rad4XCNN revealed a notable alignment between CNN-derived and radiomic features, indicating shared principles of extracting local characteristics. Moreover, we addressed concerns surrounding XAI through (Rad4XCNN), offering insights into its reliability, the accuracy-explainability trade-off dilemma, and the complexity of deep model explanations. By identifying radiomic features that exhibit a high correlation with CNN-derived features we were able to draw some important clinical conclusions. The correlation between these features and those observed in previous studies on breast cancer diagnosis underscores the clinical relevance and potential utility of our findings.
In conclusion, the use of Radiomics in CNN explanation represents a promising advancement in the development of XCDSSs. It provides a new perspective on XAI methods that deviate from the methods based on visualization maps. In addition, it provides a form of global explanation, an aspect often overlooked in current research on explainable methods in medical imaging.
@@
This research was co-funded by the Italian Complementary National Plan PNC-I.1 "Research initiatives for innovative technologies and pathways in the health and welfare sector” D.D. 931 of 06/06/2022, "DARE - DigitAl lifelong pRevEntion" initiative, code PNC0000002, CUP: B53C22006460001; and Finanziato dall’Unione europea – Next Generation EU - Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) 2022, Prot. 2022ENK9LS. Project: "EXEGETE: Explainable Generative Deep Learning Methods for Medical Image and Signal Processing", CUP: B53D23013040006.

### v2

@@ 6 Conclusion
Our findings highlight significant differences between radiomic, CNN-derived, and ViT-derived features in terms of accuracy and explainability. While deep architectures, particularly CNNs, demonstrate superior performance in terms of accuracy, radiomic features offer clearer and more comparable explanations due to their well-defined significance. Rad4XCNN revealed a notable alignment between CNN-derived and radiomic features, indicating shared principles of extracting local characteristics. Moreover, we addressed concerns surrounding XAI through Rad4XCNN, offering insights into its reliability, the accuracy-explainability trade-off dilemma, and the complexity of deep model explanations. By identifying radiomic features that exhibit a high correlation with CNN-derived features we were able to draw some important clinical conclusions. The correlation between these features and those observed in previous studies on breast cancer diagnosis underscores the clinical relevance and potential utility of our findings.
A promising future research direction would be exploring how this approach could be extended to explain features derived from ViTs, given the different feature extraction processes compared to CNNs. Additionally, incorporating structured semantic information, such as BI-RADS attributes (e.g., density, margin shape, microcalcifications, etc.), could further enrich model interpretability. These attributes could serve dual purposes: i) improving classification accuracy and ii) enabling the association of global intelligible features with extracted deep features. This exploration could open up new pathways for aligning deep learning features with clinically recognized descriptors, facilitating more meaningful insights in medical image analysis.
Summarizing, the use of Radiomics in CNN explanation represents a promising advancement in the development of XCDSSs. It provides a new perspective on XAI methods that deviate from the methods based on visualization maps. In addition, it provides a form of global explanation, an aspect often overlooked in current research on explainable methods in medical imaging.
Ethics Statement
The study was approved by the Ethical Committee of the Policlinico University Hospital ‘P. Giaccone’ of Palermo (minute N. 02/2019 of 18/02/2019). Informed consent was waived because of the retrospective nature of the study and the analysis. The data of the patients involved were processed according to the institutional privacy regulations of the Policlinico University Hospital ‘P. Giaccone’ of Palermo. The privacy rights of the human subjects involved in the study are respected because anonymized clinical and imaging data were used.
@@
This research was i) funded by the Italian Complementary National Plan PNC-I.1 "Research initiatives for innovative technologies and pathways in the health and welfare sector” D.D. 931 of 06/06/2022, "DARE - DigitAl lifelong pRevEntion" initiative, code PNC0000002, CUP: B53C22006460001; and ii) funded by the European Union – Next Generation EU - Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) 2022, Prot. 2022ENK9LS. Project: "EXEGETE: Explainable Generative Deep Learning Methods for Medical Image and Signal Processing", CUP: B53D23013040006.